Chair: Jay

Next meeting chair: Eloise

Attended: Simin, Earl, Lim, Warren, Jasper, Prad, Eloise, Jamie, Katie


General Issues

- fs updates on Hb and Ke antennas. New setup procedures. AuScope wiki notes are good to use atm, Jamie is in process of keep updating them (checklist etc.)

- Hobart control room camera flash player issue. new camera has arrived and is in the process of integrate

- slogit script isn't working properly for Hobart26 atm. Jamie will look into it Fixed - Script failed due to missing shared SSH key. Not an issue previously, but new script works by “pushing” the files from the FS machine rather than aving them “pulled”

- IP switches for Hb & ke DBBC3? ( for Ke, UPS reboot instructions for DBBC3HB)

- 'TCP/IP connection was closed by remote peer', antenna communication error on Ke. Use antenna=open running each 30s to keep communication open. Some network issue at ke, updates are happening as we speak.


Hobart 26m

  • Drives broke. Tried to reset drives few times but still no success. Tripped limit switch mechanical fault stopped operation. Fixed two limit switches. Found hydraulic leak needs fixing very soon.
  • 2300 UT schedule halted something is still bad. suspect main x-bearing in east caused limit switch fault (w)

Katherine 12m

  • DBBC3 power levels were bad, had to power-cycle it. still bad even after power cycling (Jay). DBBC3 still have some issues, Jamie's working on it to fix these. power levels for DBBC3 at ke are generally not good atm. power level 'not ok' can be ignored at the time.


Hobart 26m

  • setcl/fmset reports 1 second offset after DBBC pps_delays have drifted. Schedule halted to investigate & resynch (JMc). proc file issue.
  • large delay difference between formatter and maser delay alarms now and then. Everything is ok though. so these alarms are quite annoying (Jay). Jamie will check it out, check fmset. Error is due to missing values - it looks like clkoff/maserdelay tried to read a file while it was being updated, with a blank entry as a result. The Log Monitor should be updated to ignore these.


Ke 12m

  • FS frozen after first scan. looked closer it seems that the problem is because midob tried to call bin/, which no longer works. Changed it to call “samplers” instead. *Note: no more dbbc.log file, Jamie may want to have better fix? (Lim). new updates are working ok

Yg 12m

  • E-stop triggered again (JMc), due to bad weather. if antenna isn't switching on, have a look at timeke.


Hobart 26m

  • S/X receiver stuck in Y focus. focus platform Y drive controller had tripped over current. Reset controller, moved to < -45.0 south and got receiver on axis (warren). Hardware issue, repair parts have been ordered. for the moment if you need a receiver change at ho26, give a call to person in charge at observatory.